# Explicit $7$-torsion in the Tate–Shafarevich groups of genus $2$ Jacobians

### v1

@@ 4.2. When $\mathcal{X}$ has no rational points
We now prove Theorem 1.1. In order to apply visibility we must first show that the pairs $(C,E)$ in Table 1 are in fact $(7,\mathfrak{p})$-congruent (not simply up to quadratic twist, as we proved in Section 4).
Lemma 5.1.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $J/\mathbb{Q}$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$. Suppose that $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ and that there exists a squarefree integer $d\in\mathbb{Z}$ such that $E^{d}$ and $J$ are $(p,\mathfrak{p})$-congruent. Then $d$ is supported on the set of primes consisting of $p$, the bad primes of $E$, and the bad prime of $J$.
@@
Proposition 5.2.
For each pair $(E,C)$ of elliptic curve $E/\mathbb{Q}$ and genus 2 curve $C/\mathbb{Q}$ in Table 1 we have a $(7,\mathfrak{p})$-congruence between $E$ and $J=\operatorname{Jac}(C)$ for some choice of $\mathfrak{p}|7$ in $\mathcal{O}_{D}$.
@@
Let $\ell\neq 7$ be a good prime for $C$ and $E$. By [Fis16, (5.2)] (which follows from [FLS+01, Section 2.1] or [MS93, Lemma 3]) a $(7,\mathfrak{p})$-congruence between $E^{d}/\mathbb{Q}$ and $J/\mathbb{Q}$ gives a congruence modulo $7$
@@
Testing (2) on the divisors $d$ of the product $7$ and the bad primes of $E$ and $C$ shows that $E^{d}$ and $J$ are not $(7,\mathfrak{p})$-congruent for any $d\neq 1$ (by Lemma 5.1). By 4.4 $E$ and $J$ are $(7,\mathfrak{p})$-congruent up to quadratic twist since $E$ corresponds to a point on one of the twists $X_{J[\mathfrak{p}]}^{\pm}(7)$. It therefore follows that $E$ and $J$ are $(7,\mathfrak{p})$-congruent for some choice of $\mathfrak{p}$ dividing $7$. ∎
Using the congruences supplied by 5.2 we now prove Theorems 1.1 and 1.5 by applying [Fis16, Theorem 2.2].
Proof of Theorems 1.1 and 1.5.
@@
Let $E/\mathbb{Q}$ and $J/\mathbb{Q}$ be one of the pairs of elliptic curve and genus $2$ Jacobian from Theorem 1.1 or 1.5. We check that in each case $J$ is geometrically simple by applying the condition in [CF96, Section 14.4] and [Sto95]. The $7$-torsion subgroups of $E(\mathbb{Q})$ and $J(\mathbb{Q})$ are trivial. The rank of $E/\mathbb{Q}$ is $2$ and the rank of $J/\mathbb{Q}$ is 00 (the rank of $J/\mathbb{Q}$ is bounded using $2$-descent, which is implemented as RankBounds in Magma). For each discriminant $D$ appearing in Theorems 1.1 and 1.5 the prime $7$ not only splits in $\mathcal{O}_{D}$, but $7=\operatorname{Nm}\eta$ for some $\eta\in\mathcal{O}_{D}$. In particular the isogeny $\mathfrak{p}$ is equal to the multiplication-by-$\eta$-map on $J$, and $J(\mathbb{Q})/\mathfrak{p}J(\mathbb{Q})=0$.
By 5.2 the elliptic curve $E$ is $(7,\mathfrak{p})$-congruent to $J$. The abelian varieties $E$ and $J$ have good reduction at $7$, so by [Fis16, Theorem 2.2] it suffices to show that the Tamagawa numbers of $E/\mathbb{Q}$ and $J/\mathbb{Q}$ are coprime to $7$.
We compute the Tamagawa numbers of $E/\mathbb{Q}$ using Magma. Except for the Jacobian of conductor $3200^{2}$ in Table 1, for each bad prime $p$ of $J$ one may check that the order of the geometric component group of $J$ at $p$ is coprime to $7$ using Liu’s genus2reduction in SageMath and Donnelly’s Magma functions RegularModel and ComponentGroup.
For the Jacobian of conductor $3200^{2}$ in Table 1 the computation of the Tamagawa number of $J/\mathbb{Q}$ at $2$ was carried out in the appendix to [KS23] (where it is shown that the Tamagawa number is $1$). ∎
6. Evidence towards 1.6
@@
The Jacobian $J/\mathbb{Q}$ of $C$ has RM by $\mathbb{Z}[\sqrt{2}]$. In [KS23, A.3] it is noted that the Birch and Swinnerton-Dyer conjecture predicts $|\Sha(J^{-11}/\mathbb{Q})|=7^{2}$. By 4.4, for some choice of factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathbb{Z}[\sqrt{2}]$ we have models
@@
We were unable to find rational points on any of these curves, except on $X_{J[\mathfrak{p}]}^{\mp}(7)$ where we find exactly one point which corresponds to the elliptic curve $E/\mathbb{Q}$ with LMFDB label 1242.m1 and Weierstrass equation $y^{2}+xy+y=x^{3}-x^{2}-1666739x-2448131309$. Using the argument in 5.2 it can be shown that $E$ and $J$ are $(7,\mathfrak{p})$-congruent. However, the quadratic twist of $E$ by $-11$ has trivial Mordell-Weil group, so cannot be used to visualise the (conjectural) non-trivial elements of $\Sha(J^{-11}/\mathbb{Q})[7]$.

### v2

@@ 5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1
We now prove Theorem 1. In order to apply visibility we must first show that the pairs $(C,E)$ in Table 1.1 are in fact $(7,\mathfrak{p})$-congruent (not simply up to quadratic twist, as we proved in Section 4).
Lemma 21.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $J/\mathbb{Q}$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$. Suppose that $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ and that there exists a squarefree integer $d\in\mathbb{Z}$ such that $E^{d}$ and $J$ are $(p,\mathfrak{p})$-congruent. Then $d$ is supported on the set of primes consisting of $p$, the bad primes of $E$, and the bad primes of $J$.
@@
Proposition 22.
For each pair $(E,C)$ of elliptic curve $E/\mathbb{Q}$ and genus 2 curve $C/\mathbb{Q}$ in Table 1.1 we have a $(7,\mathfrak{p})$-congruence between $E$ and $J=\operatorname{Jac}(C)$ for some choice of $\mathfrak{p}|7$ in $\mathcal{O}_{D}$.
@@
Let $\ell\neq 7$ be a good prime for $C$ and $E$. By [Fis16, (5.2)] (which follows from [FLS+01, Section 2.1] or [MS93, Lemma 3]) a $(7,\mathfrak{p})$-congruence between $E^{d}/\mathbb{Q}$ and $J/\mathbb{Q}$ gives a congruence modulo 77
@@
Testing (2) on the divisors $d$ of the product of $7$ and the bad primes of $E$ and $C$ shows that $E^{d}$ and $J$ are not $(7,\mathfrak{p})$-congruent for any $d\neq 1$ (by Lemma 21). By 20 $E$ and $J$ are $(7,\mathfrak{p})$-congruent up to quadratic twist since $E$ corresponds to a point on one of the twists $X_{J[\mathfrak{p}]}^{\pm}(7)$. It therefore follows that $E$ and $J$ are $(7,\mathfrak{p})$-congruent for some choice of $\mathfrak{p}$ dividing $7$. ∎
Using the congruences supplied by 22 we now prove Theorems 1 and 5 by applying [Fis16, Theorem 2.2].
Proof of Theorems 1 and 5.
@@
Let $E/\mathbb{Q}$ and $J/\mathbb{Q}$ be one of the pairs of elliptic curve and genus $2$ Jacobian from Theorem 1 or 5. We check that in each case $J$ is geometrically simple by applying the condition in [CF96, Section 14.4] and [Sto95]. The $7$-torsion subgroups of $E(\mathbb{Q})$ and $J(\mathbb{Q})$ are trivial. The rank of $E/\mathbb{Q}$ is $2$ and the rank of $J/\mathbb{Q}$ is 0 (the rank of $J/\mathbb{Q}$ is bounded using $2$-descent, which is implemented as RankBounds in Magma). For each discriminant $D$ appearing in Theorems 1 and 5 the prime $7$ not only splits in $\mathcal{O}_{D}$, but $7=\operatorname{Nm}\eta$ for some $\eta\in\mathcal{O}_{D}$. In particular the isogeny $\mathfrak{p}$ is equal to the multiplication-by-$\eta$-map on $J$, and $J(\mathbb{Q})/\mathfrak{p}J(\mathbb{Q})=0$.
By 22 the elliptic curve $E$ is $(7,\mathfrak{p})$-congruent to $J$. The abelian varieties $E$ and $J$ have good reduction at $7$, so by [Fis16, Theorem 2.2] it suffices to show that the Tamagawa numbers of $E/\mathbb{Q}$ and $J/\mathbb{Q}$ are coprime to $7$.
We compute the Tamagawa numbers of $E/\mathbb{Q}$ using Magma. Except for the Jacobian of conductor $3200^{2}$ in Table 1.1, for each bad prime $p$ of $J$ one may check that the order of the geometric component group of $J$ at $p$ is coprime to $7$ using Liu’s genus2reduction in SageMath and Donnelly’s Magma functions RegularModel and ComponentGroup.
For the Jacobian of conductor $3200^{2}$ in Table 1.1 the computation of the Tamagawa number of $J/\mathbb{Q}$ at 22 was carried out in the appendix to [KS25] (where it is shown that the Tamagawa number is 11). ∎
6. Evidence towards 6
@@
The Jacobian $J/\mathbb{Q}$ of $C$ has RM by $\mathbb{Z}[\sqrt{2}]$. In [KS25, A.3] it is noted that the Birch and Swinnerton-Dyer conjecture predicts $|\Sha(J^{-11}/\mathbb{Q})|=7^{2}$. By 20, for some choice of factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathbb{Z}[\sqrt{2}]$ we have models
@@
We were unable to find rational points on any of these curves, except on $X_{J[\mathfrak{p}]}^{\mp}(7)$ where we find exactly one point which corresponds to the elliptic curve $E/\mathbb{Q}$ with LMFDB label 1242.m1 and Weierstrass equation $y^{2}+xy+y=x^{3}-x^{2}-1666739x-2448131309$. Using the argument in 22 it can be shown that $E$ and $J$ are $(7,\mathfrak{p})$-congruent. However, the quadratic twist of $E$ by 11-11 has trivial Mordell-Weil group, so cannot be used to visualise the (conjectural) non-trivial elements of $\Sha(J^{-11}/\mathbb{Q})[7]$.